OpenStudy (anonymous):

A copier was purchased by a company for $7,500. After 5 years it is estimated that the value of the copier will be $4,500. If the value in dollars V and the time the copier has been in use t are related by a linear equation, find the equation that relates V and t.

OpenStudy (anonymous):

First, find out how much the copier depreciated per year. In this case, (7500 - 4500) / 5 = 3000 / 5 = 600 per year Use this as your coefficient for annual depreciation, where t represents time, in years, so that V is the current value. V = 7500 - 600t

OpenStudy (anonymous):

V is the value of the copier in t years.